OpenStudy (anonymous):

You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly?

jimthompson5910 (jim_thompson5910):

this is a binomial distribution problem, so you'll use the formula P(X = k) = (n C k)*(p)^(k)*(1-p)^(n-k) in this case n = 5 p =1/4 = 0.25 k = 3

OpenStudy (anonymous):

i am confused about the X and the C in the equation

jimthompson5910 (jim_thompson5910):

n C k is the combination formula, so you'll either use a calculator or this formula n C k = (n!)/(k!*(n-k)!)

jimthompson5910 (jim_thompson5910):

and don't worry about the X